# Active Sensing for Multiuser Beam Tracking with Reconfigurable Intelligent Surface

### v2

@@ V-D Performance Comparison
For the benchmarks with fixed sensing schemes, Fig. 9 shows that the ones with learned sensing vectors outperform those with random sensing vectors. This shows the performance gain of designing the RIS sensing vectors based on channel statistics. Additionally, the GNN & LSTM benchmark, which uses the same RIS reflection coefficients for both pilot acquisition and data transmission stages, performs less effectively as compared to other GNN & LSTM benchmarks that employ distinct RIS reflection coefficients for these two stages. This shows the importance of designing the RIS reflection coefficients in channel sensing and data transmission stages separately. Moreover, we remark that a $5\%$ performance improvement can be obtained from power allocation refinement.
Furthermore, Fig. 9 shows that the proposed framework maintains its performance for transmission frames beyond $20$, despite being trained with only $20$ concatenated active sensing units corresponding to $20$ frames. We also observe empirically that the proposed framework continues to sustain its performance over hundreds of frames by reusing the active sensing unit. This robust performance beyond the training range is due to the ability of the LSTM cells to effectively capture the UE mobility patterns and to consistently update channel information based on the latest received pilots.
@@
Next, we evaluate the performance of the proposed approach for different number of sensing vector $L$’s within each frame. As can be seen from Fig. 10, the proposed active sensing framework always outperforms the benchmarks with fixed sensing strategies, because it leverages past observations to actively design the RIS sensing vectors. Furthermore, it is observed that LSTM-based methods (including the proposed approach) show faster convergence as the number of sensing sub-blocks $L$ increases. This is because a larger $L$ provides the LSTM with more observations in every frame, thereby facilitating more efficient capturing of temporal channel correlations.
V-E Interpretation of the Learned Solutions
@@
$r_{\text{RIS}}(\phi_{2,1},\dots,\phi_{2,N_{G}},\theta_{3,[x,y]}^{(t,0)},\phi_{3,[x,y]}^{(t,0)},\bm{v}_{\ell}^{(t)})$ (25)
$=|(\sum_{i=1}^{N_{G}}\bm{a}_{\text{RIS}}(\theta_{2,i},\phi_{2,i}))^{\sf{H}}\operatorname{diag}(\bm{v}_{\ell}^{(t)})\bm{a}_{\text{RIS}}(\theta_{3,[x,y]}^{(t,0)},\phi_{3,[x,y]}^{(t,0)})|\ ,$
@@
$r_{\text{AP},k}(\phi,\bm{b}_{k}^{(t)})=|\bm{a}_{\text{AP}}(\phi)^{\sf{H}}\bm{b}_{k}^{(t)}|\ ,k=1,\dots,K\ ,$ (26)
@@
$\frac{|(\bm{A}^{(t,n)}_{[x,y]}\widetilde{\bm{w}}^{(t)})^{\sf{H}}\bm{b}_{k}^{(t)}|^{2}}{\sum_{j=1,j\neq k}^{K}|(\bm{A}^{(t,n)}_{[x,y]}\widetilde{\bm{w}}^{(t)})^{\sf{H}}\bm{b}_{j}^{(t)}|^{2}+\sigma_{d}^{2}}\ ,\ k=1,\dots,K\ ,$ (27)
@@
Interpretation of the Downlink Beam Alignment Strategy
@@
Interpretation of the Learned Solutions via SINR Map
@@
As illustrated in Fig. 13, the proposed framework designs $L=3$ RIS sensing vectors cooperatively. For example, in the pilot stage of the $30$-th frame, one UE (represented by the red dot) is not targeted by the first RIS sensing vector $\bm{v}_{1}^{(t)}$, but is focused by $\bm{v}_{2}^{(t)}$. Meanwhile, the weaker array response of $\bm{v}_{2}^{(t)}$ towards the green UE and $\bm{v}_{3}^{(t)}$ towards the blue UE are compensated by stronger responses from the other two sensing vectors, respectively. This cooperative sensing strategy is not seen in the initial tracking frame, but is progressively developed over time by leveraging historical channel observations. As a result, the proposed framework actively narrows the sensing range and enhances the pilot SNR in the uplink CSI acquisition stages, leading to more informative neural network inputs and improved beam tracking performance. In contrast, the GNN & LSTM benchmark employs a non-adaptive approach by using channel statistics to train fixed RIS sensing vectors, which cover nearly the entire service area for CSI acquisition. However, this broad coverage results in lower uplink SNR during CSI acquisition and, subsequently, higher pilot overhead.
@@ VI Conclusion
This paper proposes an active sensing based beam tracking approach for RIS-assisted multiuser mobile communication systems. The proposed approach can adaptively design the AP beamformers and the RIS reflection coefficients for downlink data transmission, as well as the RIS reflection coefficients for uplink channel sensing. This is accomplished by utilizing an LSTM-based RNN to summarize received pilot information into state vectors, and a GNN to optimize the mapping from state vectors to the beam tracking matrices. Numerical results demonstrate that the proposed active sensing framework produces interpretable solutions and achieves significantly higher data rates than benchmark data-driven methods with non-adaptive sensing.
(a) Array response of the downlink RIS reflection coefficients $\bm{w}^{(t)}$ obtained by the proposed framework.
(b) Array response of the downlink beamforming vectors $\bm{b}^{(t)}_{k}$’s obtained by the proposed framework.
Figure 11: Array response of the downlink RIS reflection coefficients and beamformers obtained by the proposed framework in different frames. The proposed framework is tested on a randomly generated trajectory. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.
(a) SINR (dB) map of the UE $1$.
(b) SINR (dB) map of the UE $2$.
(c) SINR (dB) map of the UE $3$.
Figure 12: SINR (dB) map around the position of mobile UEs under the proposed active sensing framework in different frames. The proposed framework is tested on a randomly generated trajectory. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.
(a) $\bm{v}^{(t)}_{1}$, actively designed by the proposed framework.
@@
(c) $\bm{v}^{(t)}_{2}$, actively designed by the proposed framework.
@@
(e) $\bm{v}^{(t)}_{3}$, actively designed by the proposed framework.
@@
Figure 13: The array response of the RIS sensing vectors obtained by the proposed active sensing framework and by the GNN & LSTM benchmark without active sensing in the pilot stage (the 0-th block) across the different frames. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.

### v3

@@ V-D Performance Evaluation
For the benchmarks with fixed sensing schemes, Fig. 9 shows that the ones with learned sensing vectors outperform those with random sensing vectors. This shows the performance gain of designing the RIS sensing vectors based on channel statistics. Additionally, the GNN & LSTM benchmark that uses the same RIS reflection coefficients for both pilot acquisition and data transmission stages performs less effectively as compared to the other GNN & LSTM benchmarks that employ distinct RIS reflection coefficients for these two stages. This shows the importance of designing the RIS reflection coefficients in the channel sensing and data transmission stages separately. We also observe that the proposed approach significantly outperforms the GNN & LSTM benchmark with GNN directly designing the beamforming matrix $\bm{B}^{(t)}$ based on pilots from the first $L$ sub-blocks of each frame. This advantage arises from leveraging the beamforming structure during the training stage, enabling more efficient training and better coordination between the downlink AP beamformers and the RIS reflection coefficients. Moreover, we remark that a $5\%$ performance improvement can be obtained from power allocation refinement.
[…]
Figure 10: The impact of different numbers of sensing vectors $L$ within each frame on different beam tracking methods. Here, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.
@@
Next, we evaluate the performance of the proposed approach for different numbers of sensing vectors $L$ within each frame. As can be seen from Fig. 10, the proposed active sensing framework always outperforms the benchmarks with fixed sensing strategies, because it leverages past observations to actively design the RIS sensing vectors. Furthermore, it is observed that LSTM-based methods (including the proposed approach) show faster convergence as the number of sensing sub-blocks $L$ increases. This is because a larger $L$ provides the LSTM with more observations in every frame, thereby facilitating more efficient capturing of temporal channel correlations.
Figure 11: The impact of different numbers of AP antennas $M$ and RIS elements $N_{r}$ on the performance of the proposed approach. Here, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.
Impact of Different $M$ and $N_{r}$
In this scenario, we fix the number of UEs at $K=3$ and evaluate the performance of the proposed approach with varying numbers of AP antennas $M$ and RIS elements $N_{r}$. As shown in Fig. 11, the system performance (as measured by the minimum user rate) improves as the number of AP antennas and the number of RIS elements increase. Additionally, the proposed active sensing framework enhances tracking performance by up to $10\%$ across different configurations of $M$ and $N_{r}$. Moreover, the proposed approach consistently outperforms the GNN benchmark, which uses fixed sensing vectors learned from channel statistics.
Figure 12: Average minimum user rate versus block number for the proposed active sensing framework. Here, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.
Effectiveness of Fixing Beamforming Strategy Across the Frame
Here, we illustrate the effect of using a fixed downlink beamforming strategy within each frame. In Fig. 12, the dashed line shows the performance of the proposed framework in term of the minimum user data rate using fixed downlink beamforming vectors and RIS reflection coefficients, as the channels vary from the first block to the subsequent blocks within the frame. We see that the minimum data rate slightly decreases by about $3.5\%$ due to reusing the downlink beamforming vectors and RIS reflection coefficients for $N=3$ blocks within each data transmission stage, as the channels slowly evolve over time. In practice, the number of data transmission blocks per frame $N$ can be adjusted based on the environment’s dynamics, such as the speed of the mobile UEs, to further optimize the balance between pilot overhead and performance.
V-E Scalability and Complexity Analysis
In terms of computational complexity, it is important to note that the proposed learning-based approach is trained offline, so the main concern is inference complexity. The inference phase mainly involves matrix-vector multiplications, with the complexity scaling per LSTM cell per transmission frame is $\mathcal{O}(d^{2}+df)$ where $d$ is the size of the hidden state and cell state, and $f$ is the dimension of the vectorized input $\bar{\bm{Y}}_{k}^{(t)}$. Given that $f$ is typically much smaller than $d$ and there are $K$ LSTM cells corresponding to $K$ mobile UEs, the total complexity for the LSTM cells per transmission frame is $\mathcal{O}(Kd^{2})$ after discarding the lower order terms. On the other hand, the complexity of the GNN is $\mathcal{O}(Kd\eta+(K+N_{r}L)\kappa+(K+2)^{2}\varsigma D)$, where $D$, $\eta$, $\kappa$, and $\varsigma$ denote the number of updating layers in the GNN, the dimension of the input layer, the dimension of the output layer, and the computational complexity of the aggregation and combination operations, respectively. Therefore, the total computational complexity for the proposed active sensing unit during the inference stage is $\mathcal{O}(Kd^{2}+Kd\eta+(K+N_{r}L)\kappa+(K+2)^{2}\varsigma D)$. It is important to note that the inference process is highly parallelizable using modern graphic process units (GPUs), so that in practice the proposed approach can be executed very efficiently.
We note that there are several ways to further reduce the computational complexity of the proposed approach in practical implementations, particularly for large-scale networks with many UEs $K$ and RIS elements $N_{r}$. First, the $K$ copies of the LSTM cells can be executed in parallel, making the actual complexity equivalent to the runtime of a single LSTM cell. Second, as shown in Fig. 10, the proposed active sensing framework can achieve satisfactory performance with very few sensing vectors. Thus, only a few RIS sensing vectors need to be designed, thereby conserving computational resources. Finally, while we opt for relatively large neural networks in the simulations to characterize ultimate performance, we find that reducing the size of the neural network does not significantly impact performance.
V-F Interpretation of Learned Solutions
@@
$r_{\text{RIS}}\left(\phi_{2,1},\dots,\phi_{2,N_{G}},\theta_{3,[x,y]}^{(t,0)},\phi_{3,[x,y]}^{(t,0)},\bm{v}_{\ell}^{(t)}\right)$ (25)
$\!=\!\left|\left(\sum_{i=1}^{N_{G}}\bm{a}_{\text{RIS}}\left(\theta_{2,i},\phi_{2,i}\right)\!\right)^{\sf{H}}\!\!\!\!\operatorname{diag}\left(\bm{v}_{\ell}^{(t)}\right)\bm{a}_{\text{RIS}}\!\left(\theta_{3,[x,y]}^{(t,0)},\phi_{3,[x,y]}^{(t,0)}\right)\right|,$
@@
$r_{\text{AP},k}(\phi,\bm{b}_{k}^{(t)})=\left|\bm{a}_{\text{AP}}(\phi)^{\sf{H}}\bm{b}_{k}^{(t)}\right|,k=1,\dots,K,$ (26)
@@
$\frac{\left|\left(\bm{A}^{(t,n)}_{[x,y]}\widetilde{\bm{w}}^{(t)}\right)^{\sf{H}}\bm{b}_{k}^{(t)}\right|^{2}}{\sum_{j=1,j\neq k}^{K}\left|\left(\bm{A}^{(t,n)}_{[x,y]}\widetilde{\bm{w}}^{(t)}\right)^{\sf{H}}\bm{b}_{j}^{(t)}\right|^{2}+\sigma_{d}^{2}},\ k=1,\dots,K,$ (27)
@@
(a) Array responses of the downlink RIS reflection coefficients $\bm{w}^{(t)}$ obtained by the proposed framework.
(b) Array responses of the downlink beamforming vectors $\bm{b}^{(t)}_{k}$ obtained by the proposed framework.
Figure 13: Array responses of the downlink RIS reflection coefficients and beamformers obtained by the proposed framework in different frames. The proposed framework is tested on a randomly generated trajectory. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.
(a) SINR (dB) maps of the UE $1$.
(b) SINR (dB) maps of the UE $2$.
(c) SINR (dB) maps of the UE $3$.
Figure 14: SINR (dB) maps around the positions of mobile UEs under the proposed active sensing framework in different frames. The proposed framework is tested on a randomly generated trajectory. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.
Interpretation of Downlink Beam Alignment Strategy
@@
Interpretation of Learned Solutions via SINR Map
@@
As illustrated in Fig. 15, the proposed framework designs $L=3$ RIS sensing vectors cooperatively. For example, in the pilot stage of the $30$-th frame, one UE (represented by the red dot) is not targeted by the first RIS sensing vector $\bm{v}_{1}^{(t)}$, but is focused on by $\bm{v}_{2}^{(t)}$. Meanwhile, the weaker array response of $\bm{v}_{2}^{(t)}$ towards the green UE and $\bm{v}_{3}^{(t)}$ towards the blue UE are compensated by stronger responses from the other two sensing vectors, respectively. This cooperative sensing strategy is not seen in the initial tracking frame, but is progressively developed over time by leveraging historical channel observations. As a result, the proposed framework actively narrows the sensing range and enhances the pilot SNR in the uplink CSI acquisition stages, leading to more informative neural network inputs and improved beam tracking performance. In contrast, the GNN & LSTM Without Active Sensing benchmark employs a non-adaptive approach by using channel statistics to train fixed RIS sensing vectors, which cover a larger area for CSI acquisition. However, without adaptive focusing, this approach risks losing track of certain UEs; for instance, the green UE is not as well focused on in the $30$-th frame as compared to the proposed active sensing strategy. The broader coverage would in general result in lower uplink SNR during CSI acquisition and, subsequently, higher pilot overhead.
@@ VI Conclusion
This paper proposes an active sensing based beam tracking approach for RIS-assisted multiuser mobile communication systems. The proposed approach can adaptively design the AP beamformers and the RIS reflection coefficients for downlink data transmission, as well as the RIS reflection coefficients for uplink channel sensing. This is accomplished by utilizing an LSTM-based RNN to summarize received pilot information into state vectors, and a GNN to optimize the mapping from the state vectors to the beam tracking matrices. Numerical results demonstrate that the proposed active sensing framework produces interpretable solutions and achieves significantly higher data rates than benchmark data-driven methods with non-adaptive sensing.
(a) $\bm{v}^{(t)}_{1}$ actively designed by the proposed framework.
@@
(c) $\bm{v}^{(t)}_{2}$ actively designed by the proposed framework.
@@
(e) $\bm{v}^{(t)}_{3}$ actively designed by the proposed framework.
@@
Figure 15: The array responses of the RIS sensing vectors obtained by the proposed framework (left) and by the GNN & LSTM Without Active Sensing benchmark (right) in the pilot stage (the 0-th block) across different frames. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$, $P_{u}=5\text{dBm}$, and $P_{d}=10\text{dBm}$.